An alternating emf is given by e=100sin 314t,the time within which the emf will be half of its maximum value is
 

Grade:12

1 Answers

Arun
5 years ago
Dear Khan
 
for half emf
 
sin 314t = ½
 
sin 100 * pi * t = ½
 
100 * pi * t = pi / 6
 
t = 1/600 sec
 
Hope it helps
Regards
Arun